# Communicating Plans, Not Percepts: Scalable Multi-Agent Coordination with Embodied World Models

### v2

@@ Abstract
Robust coordination is critical for effective decision-making in multi-agent systems, especially under partial observability. A central question in Multi-Agent Reinforcement Learning (MARL) is whether to engineer communication protocols or learn them end-to-end. We investigate this dichotomy using embodied world models. We propose and compare two communication strategies for a cooperative task-allocation problem. The first, Learned Direct Communication (LDC), learns a protocol end-to-end, with agents generating messages and actions concurrently. The second, Intention Communication, uses an engineered inductive bias: a compact, learned world model, the Imagined Trajectory Generation Module (ITGM), to simulate future states. Agents then communicate a summary of this plan. We evaluate these approaches on goal-directed interaction in a grid world, a canonical abstraction for embodied AI problems. Our experiments reveal that while emergent communication is viable in simple settings, the engineered, world model-based approach shows superior performance, sample efficiency, and scalability as complexity increases. These findings advocate for integrating structured, predictive models into MARL agents to enable active, goal-driven coordination.
@@ 3.3 Training Framework
We use an on-policy REINFORCE actor-critic algorithm with a learned baseline (Sutton and Barto, 2018; NIPS1999_6449f44a). The actor and critic are separate neural networks with shared initial layers. Updates are performed at the end of each episode. We use a linear learning rate decay schedule to balance exploration and exploitation. The learning rate $\text{lr}_{e}$ for a given episode $e$ is calculated as:
@@ 4.1 Architecture and Information Flow
In LDC, a single neural network governs each agent’s decision-making, acting as a combined policy and message generator. At each timestep $t$, agent $i$’s network takes its history-stacked observation $o_{t}^{(i)}$ and the previous message from its teammate $m_{t-1}^{(j)}$ as input. The network simultaneously produces two outputs: (1) A probability distribution over the action space, from which $a_{t}^{(i)}$ is sampled. (2) A probability distribution over the message space, from which $m_{t}^{(i)}$ is sampled.
This message $m_{t}^{(i)}$ is broadcast to agent $j$ for use at timestep $t+1$. The system is trained with the shared reward signal. Gradients from the actor-critic loss flow back through both pathways, allowing agents to jointly learn to act and communicate to maximize team performance. We use a minimal binary message space $\{0,1\}$, forcing a highly compressed communication scheme.
Figure 2: Information flow in the Learned Direct Communication (LDC) model for two agents. Each agent’s network processes local observations and the teammate’s previous message to generate a new action and message. The critic network, which shares input layers, is not shown.
@@ 5 Approach 2: Intention Communication via Latent World Models
To overcome the scalability and sample-efficiency limits of emergent protocols, we designed Intention Communication. This approach is based on the principle that coordination is improved by agents sharing their future plans. We do this by embedding a compact, learned world model in each agent, providing an inductive bias for forward-looking interaction. The architecture has two components: the Imagined Trajectory Generation Module (ITGM) and the Message Generation Network (MGN).
@@ 5.1 Imagined Trajectory Generation Module (ITGM): A Latent World Model
The ITGM acts as a lightweight, deterministic world model that operates in a learned latent space. Its purpose is to forecast the agent’s future latent states, conditioned on its policy, not to predict complex observations. This "mental simulation" serves as a proxy for the agent’s intention.
Initial State Encoding
At timestep $t$, the ITGM receives the agent’s observation $o_{t}^{(i)}$ and its teammate’s previous message $m_{t-1}^{(j)}$. These are concatenated and passed through a linear encoder to produce an initial latent state vector, $z_{t}^{(i)}$:
$z_{t}^{(i)}=\text{ReLU}(\mathbf{W}_{\text{enc}}[o_{t}^{(i)}\oplus m_{t-1}^{(j)}]+\mathbf{b}_{\text{enc}})$
where $\oplus$ denotes concatenation. This vector $z_{t}^{(i)}$ is a compressed representation of the agent’s current world knowledge.
Latent Rollout
The core of the ITGM is a learned latent transition model, implemented as a simple feed-forward network. This model unrolls a trajectory for a fixed horizon $H$ (e.g., $H=5$). For each step $k$ in the horizon ($k=0,\dots,H-1$):
$z_{t+k+1}^{(i)}=\text{ReLU}(\mathbf{W}_{\text{fwd}}z_{t+k}^{(i)}+\mathbf{b}_{\text{fwd}})$
This iterative process generates a sequence of predicted future latent states, $\tau^{(i)}=\langle z_{t+1}^{(i)},z_{t+2}^{(i)},\dots,z_{t+H}^{(i)}\rangle$, which constitutes the imagined trajectory.
This simple, feed-forward design was chosen for computational efficiency, a critical factor given our resource constraints. It provides an effective way for the agent to reason about its policy’s immediate consequences, forming a basis for communication.
@@ 5.2 Message Generation Network (MGN): Summarizing Intentions
The imagined trajectory $\tau^{(i)}$ is a high-dimensional sequence, unsuitable for direct transmission as a message. The MGN distills this trajectory into a compact, fixed-length message vector $m_{t}^{(i)}$ that summarizes the agent’s plan. We employ a multi-head self-attention mechanism (Vaswani et al., 2017) for this purpose. Self-attention is well-suited for this, as it lets the network learn the relative importance of points in the future trajectory. For example, the final state in the rollout might be more critical to communicate than intermediate states.
The sequence $\tau^{(i)}$ serves as the query, key, and value inputs for the attention mechanism:
$\mathbf{Q}=\tau^{(i)}\mathbf{W}_{Q},\quad\mathbf{K}=\tau^{(i)}\mathbf{W}_{K},\quad\mathbf{V}=\tau^{(i)}\mathbf{W}_{V}$
The attention output is a re-weighted trajectory sequence, highlighting salient features. This output is then mean-pooled across time to produce the final message vector $m_{t}^{(i)}$. This structured message is more informative than LDC’s single bit, as it is grounded in a forward-looking simulation.
5.3 Integrated Plan-Then-Act Framework
The ITGM and MGN modules are integrated into a "plan-then-act" cycle for each agent $i$ at each timestep $t$:
@@ 5.3 Integrated Plan-Then-Act Framework
    Receive: The agent receives its local observation $o_{t}^{(i)}$ and its teammate’s previous intention message, $m_{t-1}^{(j)}$.
@@
    Plan (Imagine): The ITGM uses these inputs to generate the imagined future trajectory $\tau^{(i)}$.
@@
    Summarize (Encode Message): The MGN processes $\tau^{(i)}$ to produce the agent’s own intention message for the current step, $m_{t}^{(i)}$. This message is broadcast to its teammate for use at step $t+1$.
@@
    Act: The agent’s policy (actor) and value (critic) networks decide based on the environmental observation and its own new intention. The policy input is $o_{t}^{(i)}\oplus m_{t}^{(i)}]$.
    $a_{t}^{(i)},V(s_{t})\sim\pi_{\theta_{i}}(\cdot|o_{t}^{(i)}\oplus m_{t}^{(i)})$
This architecture creates a powerful feedback loop. An agent’s action is directly conditioned on the plan it has just formulated and communicated, which enforces consistency between communication and behavior. The entire model (ITGM, MGN, actor-critic) is trained end-to-end. Gradients from the task reward flow through all components, allowing the agent to simultaneously learn what plans lead to high rewards, how to summarize them into messages, and how to interpret its teammate’s messages to refine its actions.
@@ 7 Conclusion and Future Work
We thank Denon Chong Cheng Zong and Jeff Lee for their contributions to research on MARL communication. We thank Dr. Matthew Berland for their mentorship and feedback.

### v3

@@ Abstract
Robust coordination is critical for effective decision-making in multi-agent systems, especially under partial observability. A central question in Multi-Agent Reinforcement Learning (MARL) is whether to engineer communication protocols or learn them end-to-end. We investigate this dichotomy using embodied world models. We propose and compare two communication strategies for a cooperative task-allocation problem. The first, Learned Direct Communication (LDC), learns a protocol end-to-end. The second, Intention Communication, uses an engineered inductive bias: a compact, learned world model, the Imagined Trajectory Generation Module (ITGM), which uses the agent’s own policy to simulate future states. A Message Generation Network (MGN) then compresses this plan into a message. We evaluate these approaches on goal-directed interaction in a grid world, a canonical abstraction for embodied AI problems, while scaling environmental complexity. Our experiments reveal that while emergent communication is viable in simple settings, the engineered, world model-based approach shows superior performance, sample efficiency, and scalability as complexity increases. These findings advocate for integrating structured, predictive models into MARL agents to enable active, goal-driven coordination.
@@ 3.3 Training Framework
We use an on-policy Advantage Actor-Critic (A2c) algorithm with a learned baseline (Sutton and Barto, 2018; NIPS1999_6449f44a). The actor and critic are separate neural networks with shared initial layers. Updates are performed at the end of each episode. We use a linear learning rate decay schedule to balance exploration and exploitation. The learning rate $\text{lr}_{e}$ for a given episode $e$ is calculated as:
@@ 4.1 Architecture and Information Flow
In LDC, each agent’s policy network $\pi_{\theta_{i}}$ outputs both an action and a message: $\pi_{\theta_{i}}(a_{t}^{(i)},m_{t}^{(i)}|o_{t}^{(i)},m_{t-1}^{(j)})$, where $j\neq i$. We keep the message space minimal to facilitate learning: a single binary value, $m_{t}^{(i)}\in\{0,1\}$. The network outputs logits for the action (passed through softmax) and message (passed through sigmoid). During training, both are sampled. The message $m_{t}^{(i)}$ is broadcast to agent $j$ for use at timestep $t+1$. The entire system is trained end-to-end, with gradients from the shared task reward shaping both the policy and the protocol. The detailed network architecture for Learned Direct Communication is depicted in figure 2.
@@ 5 Approach 2: Intention Communication via Latent World Models
To address the limitations of emergent protocols, we designed Intention Communication, an architecture with strong inductive biases for forward-looking coordination. The core idea is for agents to share explicit information about their future plans, allowing their teammates to anticipate their behavior and avoid conflicts (Liu et al., 2021). This is operationalized through a two-stage process: an agent first generates an internal mental simulation of its future trajectory, and then compresses this plan into a message for its teammate.
@@ 5.1 Imagined Trajectory Generation Module (ITGM): A Latent World Model
The ITGM serves as a compact, learned module to generate a plan. It combines a learned latent world model with the agent’s own policy to generate a short-term forecast. Its objective is to simulate a future trajectory that would result from following its current policy. This policy-conditioned simulation provides the raw material for forming an intention. The process is as follows.
Initial State Encoding:
At timestep $t$, the module first takes the agent’s current local observation $o_{t}^{(i)}$ (a $(4\times F)$-dimensional vector with $F=4$) and the received message from its teammate in the previous step, $m_{t-1}^{(j)}$. (This message $m_{t-1}^{(j)}$ is first embedded and flattened into a fixed-length vector). These are concatenated and passed through a linear encoder to produce an initial $D_{hidden}$-dimensional latent state vector, $z_{t}^{(i)}$: $z_{t}^{(i)}=\text{ReLU}(\mathbf{W}_{enc}[o_{t}^{(i)}\oplus\text{embed}(m_{t-1}^{(j)})]+\mathbf{b}_{enc})$.
where $\oplus$ denotes vector concatenation. This initial state $z_{t}^{(i)}$ represents a compressed summary of the agent’s current knowledge and serves as the starting point for the plan.
Latent Rollout:
The ITGM generates the plan by unrolling a trajectory for a fixed horizon $H$. This rollout actively uses the agent’s own policy head ($\pi_{act}$) and a learned latent transition model ($f_{trans}$). This transition model is a network (e.g., an MLP) that predicts the next latent state given the current latent state and a *planned action*.
For each step $k$ in the horizon ($k=0,\dots,H-1$), a planned action is sampled from the agent’s policy and used by the latent transition model to predict the next state. This iterative process is implemented as follows, where $\text{embed}(\cdot)$ is an action embedding layer and $\pi_{act}$ is the same policy head used to select the agent’s real action:
$\tilde{a}_{t+k}\sim\pi_{act}(\cdot|z_{t+k}^{(i)})$
$z_{t+k+1}^{(i)}=\text{ReLU}(\mathbf{W}_{trans}[z_{t+k}^{(i)}\oplus\text{embed}(\tilde{a}_{t+k})]+\mathbf{b}_{trans})$
This iterative process generates a sequence of predicted future latent states, which forms the plan (the policy-conditioned trajectory): $\tau^{(i)}=\langle z_{t+1}^{(i)},z_{t+2}^{(i)},\dots,z_{t+H}^{(i)}\rangle$.
This feed-forward design was chosen for its computational efficiency, which is critical given our resource constraints. Both the encoder, the transition model, and the forward model are feed-forward networks with $D_{hidden}$ hidden units and ReLU activation functions.
@@ 5.2 Message Generation Network (MGN): Summarizing Intentions
The role of the MGN is to distill the high-dimensional plan (the latent trajectory $\tau^{(i)}$) into a compact, fixed-length message vector $m_{t}^{(i)}$ that effectively summarizes the agent’s intention. A raw sequence of latent states is too unwieldy to use as a direct communication signal. Instead, we use a self-attention mechanism to identify and aggregate the most salient features of the planned trajectory.
[…]
This attention output is aggregated via mean pooling to produce a single vector $v_{pooled}$. This vector is then passed through two hidden layers, $h_{1}=\text{ReLU}(\mathbf{W}_{1}v_{pooled}+b_{1})$ and $h_{2}=\text{ReLU}(\mathbf{W}_{2}h_{1}+b_{2})$.
To produce the message policy, $h_{2}$ is fed into $N_{comp}$ separate linear heads. Each head $k\in[1,N_{comp}]$ outputs a vector of $N_{sym}$ logits: $\text{logits}^{(k)}=\mathbf{W}_{head}^{(k)}h_{2}+b_{head}^{(k)}$.
This defines $N_{comp}$ independent categorical distributions, $\pi_{msg}^{(k)}=\text{Softmax}(\text{logits}^{(k)})$, one for each component of the message. This structured, multi-part discrete message is far more expressive than the single bit used in LDC.
5.3 Integration and End-to-End Training
The ITGM and MGN modules are seamlessly integrated into the agent’s decision-making process. The full forward pass for a single agent $i$ at timestep $t$ explicitly separates the action and message generation to respect the constraint:
@@ 5.3 Integration and End-to-End Training
    The agent receives its observation $o_{t}^{(i)}$ and the previous message from its teammate, $m_{t-1}^{(j)}$.
@@
    These inputs are encoded to produce the initial latent state $z_{t}^{(i)}$.
@@
    (Action Branch) The agent’s policy head computes the real action distribution using this initial state: $\pi_{act}(a_{t}^{(i)}|z_{t}^{(i)})$. An action $a_{t}^{(i)}$ is sampled.
@@
    (Plan/Message Branch) In parallel, the ITGM uses $z_{t}^{(i)}$ as its starting point to generate the plan $\tau^{(i)}=\langle z_{t+1}^{(i)},\dots,z_{t+H}^{(i)}\rangle$. This rollout process re-uses the same policy head $\pi_{act}$ to sample hypothetical future actions $\tilde{a}_{t+k}$ at each step of the simulation.
  5. 5.
    The MGN processes the resulting plan $\tau^{(i)}$ to output the $N_{comp}\times N_{sym}$ logits for the $N_{comp}$ message distributions $\pi_{msg}^{(k)}$.
  6. 6.
    A message $m_{t}^{(i)}=\langle m_{t}^{(1)},\dots,m_{t}^{(N_{comp})}\rangle$ is sampled by taking one symbol from each of the $N_{comp}$ categorical distributions. This composite message vector is broadcast to its teammate for use at step $t+1$.
  7. 7.
    The entire model is trained end-to-end using the A2C loss function. Gradients from the $L_{\text{actor}}$ term (specifically the $A_{t}\sum_{k}\log\pi_{msg}^{(k)}(m_{t}^{(k)})$ component) flow back to update the weights of the MGN, its hidden layers, the ITGM’s transition model ($\mathbf{W}_{trans}$), and the initial encoder ($\mathbf{W}_{enc}$). The gradients from both the action ($A_{t}\log\pi_{act}(a_{t})$) and message terms also update the shared policy head $\pi_{act}$.
The entire model is trained end-to-end using the A2C loss function, allowing gradients to flow back through all modules.
@@ 7 Conclusion and Future Work
We thank Denon Chong Cheng Zong and Jeff Lee for their contributions to background research on MARL communication. We thank Dr. Matthew Berland for their mentorship and feedback.
@@ A.2 Conclusions from Additional Experiments
A.3 Key Training and Architectural Hyperparameters
We share the hyperparameters used in this experiment for greater reproducability in table 7.
Table 7: Key Training and Architectural Hyperparameters
Parameter Value
Training Hyperparameters
Initial LR ($\text{lr}_{0}$) $1\times 10^{-3}$
Final LR ($\text{lr}_{f}$) $1\times 10^{-4}$
Discount Factor ($\gamma$) 0.99
Optimizer Adam [Kingma and Ba, 2014]
Critic Loss Coefficient ($\alpha_{c}$) 0.5
Max Episode Length ($T_{max}$) 200
Architectural Hyperparameters (all)
Frame Stack ($F$) 4
Hidden Dimension ($D_{hidden}$) 64
Hidden Layers ($N_{hidden}$) 2
Message Loss Coefficient ($\beta$) 0.1
Architectural Hyperparameters (Intention Communication)
ITGM Rollout Horizon ($H$) 3
Attention Heads ($N_{heads}$) 1
Message Components ($N_{comp}$) 8
Symbol Vocabulary Size ($N_{sym}$) 8
A.4 LDC Architecture
The detailed network architecture for Learned Direct Communication is depicted in figure 2.
Figure 2: The LDC architecture. Each agent’s policy network takes its local observation and the incoming message from the previous timestep to produce an action and an outgoing message for the next timestep. The critic network estimates the state value.